# Exploiting Symmetry in Dynamics for Model-Based Reinforcement Learning with Asymmetric Rewards

### v1

@@ IV Experiments
In this section, we demonstrate the performance of this method on two numerical examples involving learning the dynamics of a system. Two environments that we used were “Parking” from “Highway-env” [highway-env] and “Reacher” from OpenAI Gym [brockman2016openai] exhibiting rotation and/or translation symmetry which are explained in Figure 2. In “Parking” there are two controlled vehicles that we need to park to corresponding parking spots and in “Reacher” we control a robot arm with 2 joints to reach to a target position. To be able to compare our method fairly to a baseline we generated a static dataset of transitions ${(x,u,x^{\prime})}$, and we learn a dynamical model from this dataset. Consequently, the comparison between the two methodologies was centered on their respective capabilities in dynamics learning rather than policy generation. This approach ensures a fair comparison, as it relies on the same offline dataset for both methods, avoiding discrepancies that may arise from online training and dataset variations.
The dataset is constructed by training an agent using Soft Actor-Critic [haarnoja_soft_2018], a model-free reinforcement learning method. We use D3rlpy [d3rlpy] to train dynamical models from the pre-collected dataset. We implement the symmetry method discussed above using (5), where $\bar{F}$ is the neural network whose parameters are trained. The observation error (on the test dataset) of this dynamical model is compared against the observation error of learning $F$ directly. We further compare these methods (with and without symmetry) across a variety of parameter sizes. The rollouts from the resulting policy, the datasets used in this work and the code to reproduce the results in this paper can be found on our GitHub repository [^1] https://github.com/YasinSonmez/symmetry-cs285.
@@ IV-A Two Cars Parking Scenario
We use the parking environment from [highway-env] with two cars with separate goal positions. The aim of this environment is to park the two cars at their goal parking spots by controlling them separately and not allowing them to crash into obstacles or to each other.
Figure 2: Experimental environments included: (1) “Reacher” on the left featuring two rotating controlled joints that exhibit rotation symmetry with the objective of reaching a target point, and (2) “Parking” involving two controlled vehicles maneuvering to park in designated spots without collision, demonstrating both rotational and translational symmetry in dynamics.
Note that the dynamics of each car are translation- and rotation-invariant, but the reward function (relating to the location and orientation of the parking spot) does not follow the same symmetry. Therefore it is not possible to employ methods that assume that both the dynamics and the reward function are symmetric. The car dynamics in this environment follow the kinematic bicycle model [kinematicBicycle], with state $x=(y,z,v_{y},v_{z},h_{y},h_{z})$ corresponding to the $y$ position, $z$ position, $y$ component of velocity, $z$ component of velocity, cosine of heading angle, and sine of heading angle.
The state of this parking environment is 24-dimensional, with the first 6 states corresponding to the first car, the second 6 states corresponding to the second car, the third 6 states corresponding to the first car’s goal, and the final 6 states corresponding to the sixth car’s goal. This state corresponds to 4 independent systems: the two cars follow the kinematic bicycle model, and the two goal systems are static, following the simple dynamical model of $g_{k+1}=g_{k}$. We apply the method presented in Sections II and III to construct transformation groups for each of the four systems. The joint system then satisfies the transformation group formed as the product of the individual transformation groups.
@@
Figure 3: Illustration of car dynamics demonstrating translational and rotational invariance. By applying Cartan’s moving frame method, coordinates are transformed to position the car at the origin with a neutral orientation. The function $\rho$ reduces the system’s state to a lower-dimensional space, without losing essential dynamics information. Subsequently, dynamics are learned in these modified coordinates via a smaller neural network (bottom) compared to the usual NN (middle). The NN’s output is then reconverted to original coordinates using $\phi_{\gamma(x)}^{-1}$.
We first describe the transformation group applied to learn translation- and rotation-invariant dynamics for the two car models. We parameterize the symmetry group $G=SE(2)$ by a translation $(y^{\prime},z^{\prime})$ and a rotation $\theta^{\prime}$. Define the rotation matrix
$R_{\theta^{\prime}}=\begin{bmatrix}\cos\theta^{\prime}&-\sin\theta^{\prime}\\ \sin\theta^{\prime}&\cos\theta^{\prime}\end{bmatrix}.$ (7)
To define the action of $G$ on the state space, let $(y^{\prime},z^{\prime},\theta^{\prime})\in G$ and $x$ be the 6-dimensional state of the car:
$\phi_{(y^{\prime},z^{\prime},\theta^{\prime})}(x)$ $=\begin{bmatrix}R_{\theta^{\prime}}&0&0\\ 0&R_{\theta^{\prime}}&0\\ 0&0&R_{\theta^{\prime}}\end{bmatrix}x+\begin{bmatrix}\begin{bmatrix}y^{\prime}\\ z^{\prime}\end{bmatrix}\\ 0\\ 0\end{bmatrix}$ (8)
@@
In this case, the action of $G$ on the control input is the identity: $\psi_{\alpha}(u)=u$ for all $\alpha\in G$.
We now derive $\gamma(x)$, $\gamma(x)^{-1}$ (group inverse), and $\rho$ for this transformation group, so we can evaluate (5). Define $\mathcal{C}$ to be the set of states whose position is equal to 0 ($y=z=0$) and whose heading is equal to 0 ($h_{y}=1$ and $h_{z}$ = 0). The moving frame $\gamma$ is then defined as the solution to:
$\begin{bmatrix}0\\ 0\\ 1\\ 0\end{bmatrix}=\phi_{\gamma(x}^{a}(x)=\begin{bmatrix}R_{\theta^{\prime}}&0\\ 0&R_{\theta^{\prime}}\end{bmatrix}\begin{bmatrix}y\\ z\\ h_{y}\\ h_{z}\end{bmatrix}+\begin{bmatrix}y^{\prime}\\ z^{\prime}\\ 0\\ 0\end{bmatrix},$
which is given by:
We can then compute the group inverse of $\gamma(x)$ as:
Substituting $\gamma(x)$ into the definition of $\rho$, we get:
Now we present the transformation group applied to the goal dynamical model. The goal dynamical model satisfies the trivial transformation group parameterized by $(\mathbb{R}^{6},+)$ such that $\phi_{\delta}(g)=g+\delta$, and there is no input. The “a” component of $\phi$ has 6 components, leaving the “b” component with 0. Effectively, we remove the goal states entirely.
Using the above transformation groups applied to each car and goal, the transformation group applied to the joint state of the environment gives a $\rho$ which reduces the state size from 24 to 4. Thus, we learn a neural network with significantly reduced input size. Note that the computation of these functions only depends on the symmetry and not on the dynamics, so the dynamical model need not be known apriori and can be learned using the data. This procedure is explained in Figure 3.
The three neural network configurations that we used in our experiments are: 1 hidden layer with size 128, 2 hidden layers with size 128, and 3 hidden layers with size 128. In the method using symmetry, the input size is 8 (reduced coordinate size for each car of 2 and control input size of 2 for each of the two cars), and the output size is 24. In the method not using symmetry, the input size is 28 (state space size of 24 and control input size of 2 for each of the two cars), and the output size is 24.
The results of observation error vs. number of parameter updates are shown in Figure 4. At the lower number of parameters, the dynamical model training with symmetry outperforms the default method without symmetry by achieving a lower observation error. When there are two hidden layers, the method without symmetry learns slower than does the method with symmetry, but ultimately catches up in performance. When there are three hidden layers, the method without symmetry exceeds the performance of that of the method with symmetry.
@@ IV-B Reacher
In this experiment, we consider the standard reacher environment in OpenAI Gym. The dynamics of the reacher exhibit rotational symmetry with the first joint angle, but the target does not since it is fixed. We use symmetry reduction to learn model dynamics which are rotation-invariant.
The state of the reacher is 11-dimensional, with $x_{1}$ and $x_{2}$ representing the cosines of the first and second joint angles, $x_{3}$ and $x_{4}$ representing the sines of the joint angles, $x_{5}$ and $x_{6}$ representing the x- and y-coordinates of the target, $x_{7}$ and $x_{8}$ representing the angular velocities of the first and second arm, $x_{9}$ and $x_{10}$ representing the distance between the reacher fingertip and the target along the x- and y-axes, and $x_{11}$ equaling the constant 0 representing the $z$ coordinate of the fingertip.
We now describe the transformation group applied to learn a rotation-invariant dynamical model for the reacher. For $x_{5},x_{6}$, and $x_{11}$, we apply the same symmetry used in Section IV-A to describe constants. Let $G$ be parameterized by $\alpha=(\theta^{\prime},\delta_{1},\delta_{2},\delta_{3})\in\mathbb{R}^{4}$, where $\theta^{\prime}$ represents the angle by which the reacher is rotated, and $\delta$ represents translations of the target position and $x_{11}$. We define the action of $G$ on the state space as:
$\phi_{\alpha}(x)=\begin{bmatrix}\cos(\theta^{\prime})x_{1}-\sin(\theta^{\prime})x_{3}\\ x_{2}\\ \sin(\theta^{\prime})x_{1}+\cos(\theta^{\prime})x_{3}\\ x_{4}\\ \cos(\theta^{\prime})(x_{5}+\delta_{1})-\sin(\theta^{\prime})(x_{6}+\delta_{2})\\ \sin(\theta^{\prime})(x_{5}+\delta_{1})+\cos(\theta^{\prime})(x_{6}+\delta_{2})\\ x_{7}\\ x_{8}\\ \cos(\theta^{\prime})(x_{9}-\delta_{1})-\sin(\theta^{\prime})(x_{10}-\delta_{2})\\ \sin(\theta^{\prime})(x_{9}-\delta_{1})+\cos(\theta^{\prime})(x_{10}-\delta_{2})\\ x_{11}+\delta_{3}\end{bmatrix}.$ (12)
The action of $G$ on the control input is the identity: $\psi_{\alpha}(u)=u$ for all $\alpha\in G$.
To define the symmetry reduction, we define $\mathcal{C}$ by the states $x$ for which $x_{1}=1$, $x_{3}=0$, $x_{5}=0$, $x_{6}=0$, and $x_{11}=0$. The solution for the moving frame $\gamma(x)$ is
$\gamma(x)=\begin{bmatrix}\mathrm{arctan2}(-x_{3},x_{1})\\ -x_{5}\\ -x_{6}\\ -x_{11}\end{bmatrix},$
@@
Using this symmetry group, the input size of the neural network is 8 (reduced coordinate size 6 and control input size of 2), and the output size is 11. In the method not using symmetry, the input size is 13 (state space size of 11 and control input size of 2), and the output size is 11. The three neural network configurations that we used in our experiments are: 1 hidden layer with size 64, 2 hidden layers with size 64, and 3 hidden layers with size 64. We used 64 neuron size compared to 128 in “Parking” because of the simplicity of this problem.
@@ V Conclusion
This study highlights the benefits of exploiting symmetry when learning dynamical models. By focusing on the symmetry inherent in dynamics, independent of the reward function, we broaden the range of problems where symmetry can be effectively utilized. We investigated symmetries such as rotation and translation symmetries through Cartan’s method of moving frames and our proposed method proves to be effective in leveraging these symmetries to learn more accurate models, particularly at low numbers of model parameters. The experiments confirm the potential of our approach to enhance learning efficiency in MBRL settings. We compare our method with varying parameter sizes and see a higher benefit of utilizing symmetry in smaller networks.
@@ VI Acknowledgements
We thank Sergey Levine for pointing to relevant references and for discussions on the preliminary results. Additionally, we thank Hussein Sibai for discussions on symmetry reduction that eventually led to this formulation of the paper.

### v2

@@ IV Experiments
In this section, we demonstrate the performance of this method on two numerical examples involving learning the dynamics of a system. The two environments that we used were “Parking” from “Highway-env” [highway-env] and “Reacher” from OpenAI Gym [brockman2016openai]. These exhibit rotational and/or translational symmetry, as depicted in Figure 2. In “Parking” there are two controlled vehicles that we need to park to corresponding parking spots, and in “Reacher” we control a robot arm with 2 joints to reach to a target position. To be able to compare our method fairly to a baseline we generated a static dataset of transitions ${(x,u,x^{\prime})}$, and we learn a dynamical model from this dataset. Consequently, the comparison between the two methodologies was centered on their respective capabilities in dynamics learning rather than policy generation. This approach ensures a fair comparison, as it relies on the same offline dataset for both methods, avoiding discrepancies that may arise from online training and dataset variations.
The dataset is constructed by training an agent using Soft Actor-Critic [haarnoja_soft_2018], a model-free reinforcement learning method. We use D3rlpy [d3rlpy] to train dynamical models from the pre-collected dataset. We implement the symmetry method discussed above using (8), where $\bar{F}$ is the neural network whose parameters are trained. The observation error (on the test dataset) of this dynamical model is compared against the observation error of learning $F$ directly. We further compare these methods (with and without symmetry) across a variety of parameter sizes. The rollouts from the resulting policy, the datasets used in this work and the code to reproduce the results are available in a GitHub repository [^1] https://github.com/YasinSonmez/symmetry-cs285.
@@ IV-A Two Cars Parking Scenario
We use the parking environment from [highway-env] with two cars with separate goal positions. The aim is to park the two cars at their goal spots by controlling them separately and not allowing them to crash into obstacles or to each other. The dynamics of each car are translation- and rotation-invariant, but the reward function (relating to the location and orientation of the parking spot) is not. Therefore, methods that assume both the dynamics and the reward function satisfy the same symmetries are not applicable.
Figure 2: Experimental environments included: (1) “Reacher” on the left, featuring two rotating controlled joints that exhibit rotation symmetry with the objective of reaching a target point, and (2) “Parking” on the right, involving two controlled vehicles maneuvering to park in designated spots without collision, demonstrating both rotational and translational symmetry in dynamics.
The state of this parking environment is 24-dimensional, with the first 6 states corresponding to the first car, the second 6 states corresponding to the second car, the third 6 states corresponding to the first car’s goal, and the final 6 states corresponding to the second car’s goal. The state incorporates the two car models and two goal models. The goals follow the model $g_{k+1}=g_{k}$. We include the goal states in the state representation because it aligns with the conventional RL framework where all states are utilized directly in the training process without explicit distinction. We apply the method presented in Sections II and III to construct transformation groups for each of the four systems. The joint system then satisfies the transformation group formed as the product of the individual transformation groups.
@@
Figure 3: Illustration of translational and rotational invariance in car dynamics. By applying Cartan’s moving frame method, coordinates are transformed to position the car at the origin with a neutral orientation. The function $\rho$ reduces the system’s state to a lower-dimensional space, without losing essential dynamics information. The dynamics are learned in these modified coordinates via a smaller neural network (bottom) compared to the usual NN (middle). The NN’s output is then reconverted to original coordinates using $\phi_{\gamma(x)}^{-1}$. As usual, the neural network training is simplified by outputting the difference between the next and current states, represented as $\Delta F(x,u)$ and $\Delta\bar{F}(x,u)$ in Section III. This is an illustration of a single car. In the experiments, additional symmetries are observed due to the presence of two cars; however, the second car has been omitted here for clarity.
Each car has state $x=(y,z,v_{y},v_{z},h_{y},h_{z})$, corresponding to the $y$ position, $z$ position, $y$ component of velocity, $z$ component of velocity, cosine of heading angle, and sine of heading angle. We use the group $G=\mathrm{SE}(2)$ to represent the symmetries. See Example 1 for the action of $G$ on the state space and the solutions for $\gamma$ and $\rho$. The action of $G$ on the control input space is the identity: $\psi_{g}(u)=u$ for all $g\in G$. The goal dynamical model satisfies the trivial transformation group $(\mathbb{R}^{6},+)$ which acts on the goal by $\phi_{\delta}(g)=g+\delta$. There is no input to act on. The $a$ component of $\phi$ has 6 components, leaving the $b$ component with 0.
@@
Using the above transformation groups applied to each car and goal, the transformation group applied to the joint state of the environment gives a $\rho$ which transforms the input state from a 24-dimensional vector to a 4-dimensional one. Thus, we learn a neural network with significantly reduced input size. Note that the computation of these functions only depends on the symmetry and not on the dynamics. The dynamical model need not be known apriori and can be learned using the data; see Figure 3.
The three neural network configurations used in the experiments are: 1 hidden layer with size 128, 2 hidden layers with size 128, and 3 hidden layers with size 128. When using symmetry, the input size is 8 (reduced coordinate size for each car of 2 and control input size of 2 for each car), and the output size is 24. In the method not using symmetry, the input size is 28 (state space size of 24 and control input size of 2 for each of the two cars), and the output size is 24.
The results of observation error vs. number of parameter updates are shown in Figure 4. At the lower number of parameters, the dynamical model training with symmetry outperforms the default method without symmetry by achieving a lower observation error. When there are two hidden layers, the method with symmetry learns faster than the default method without symmetry, although eventually the method without symmetry catches up in performance. When there are three hidden layers, the method without symmetry exceeds the performance of that of the method with symmetry. This suggests that the relative benefit of leveraging symmetry compared to the benefit from the number of parameters diminishes as the number of neural network parameters increases. One possible cause for this is that our method of leveraging symmetry removes redundant information, but the neural network might be able to exploit this redundant information when it has a sufficiently high number of parameters.
@@ IV-B Reacher
In this experiment, we consider the standard reacher environment in OpenAI Gym. The dynamics exhibit rotational symmetry with the first joint angle, but the dynamics of the target do not since the target is fixed. We use symmetry reduction to learn model dynamics which are rotation-invariant.
The state of the reacher is 11-dimensional, with $x_{1}$ and $x_{2}$ representing the cosines of the first and second joint angles, $x_{3}$ and $x_{4}$ representing the sines of the joint angles, $x_{5}$ and $x_{6}$ representing the x- and y-coordinates of the target, $x_{7}$ and $x_{8}$ representing the angular velocities of the first and second arm, $x_{9}$ and $x_{10}$ representing the distance between the reacher fingertip and the target along the x- and y-axes, and $x_{11}$ equaling 0, representing the $z$ coordinate of the fingertip.
We now describe the transformation group applied to learn a rotation-invariant dynamical model. For $x_{5},x_{6}$, and $x_{11}$, we apply the same symmetry used in Section IV-A to describe constants. Let $G$ be parameterized by $g=(\theta^{\prime},\delta_{1},\delta_{2},\delta_{3})\in\mathbb{R}^{4}$, where $\theta^{\prime}$ represents the angle by which the reacher is rotated, and $\delta$ represents translations of the target position and $x_{11}$. We define the action of $G$ on the state space as:
$\phi_{g}(x)=\begin{bmatrix}\cos(\theta^{\prime})x_{1}-\sin(\theta^{\prime})x_{3}\\ x_{2}\\ \sin(\theta^{\prime})x_{1}+\cos(\theta^{\prime})x_{3}\\ x_{4}\\ \cos(\theta^{\prime})(x_{5}+\delta_{1})-\sin(\theta^{\prime})(x_{6}+\delta_{2})\\ \sin(\theta^{\prime})(x_{5}+\delta_{1})+\cos(\theta^{\prime})(x_{6}+\delta_{2})\\ x_{7}\\ x_{8}\\ \cos(\theta^{\prime})(x_{9}-\delta_{1})-\sin(\theta^{\prime})(x_{10}-\delta_{2})\\ \sin(\theta^{\prime})(x_{9}-\delta_{1})+\cos(\theta^{\prime})(x_{10}-\delta_{2})\\ x_{11}+\delta_{3}\end{bmatrix}.$ (10)
The action of $G$ on the control input is the identity: $\psi_{g}(u)=u$ for all $g\in G$. We define $\mathcal{C}$ by the states $x$ for which $x_{1}=1$, $x_{3}=0$, $x_{5}=0$, $x_{6}=0$, and $x_{11}=0$. The solution for the moving frame $\gamma(x)$ is
$\gamma(x)=\begin{bmatrix}\mathrm{arctan2}(-x_{3},x_{1})\\ -x_{5}\\ -x_{6}\\ -x_{11}\end{bmatrix}$
@@
Using this symmetry group, the input size of the neural network is 8 (reduced coordinate size 6 and control input size of 2), and the output size is 11. When not using symmetry, the input size is 13 (state space size of 11 and control input size of 2), and the output size is 11. The three neural network configurations that we used in our experiments are: 1 hidden layer with size 64, 2 hidden layers with size 64, and 3 hidden layers with size 64. We used 64 neuron size compared to 128 in “Parking” because of the simplicity of this problem.
@@ V Conclusion
This study highlights the benefits of exploiting symmetry when learning dynamical models. By focusing on the symmetry inherent in dynamics, independent of the reward function, we broaden the range of problems where symmetry can be effectively utilized. We investigated symmetries such as rotation and translation symmetries through Cartan’s method of moving frames and our proposed method proves to be effective in leveraging these symmetries to learn more accurate models, particularly at low numbers of model parameters. The experiments confirm the potential of our approach to enhance learning efficiency. Future work involves training policies using the proposed symmetry-enhanced method and evaluating their performance. Additionally, a theoretical framework that includes discrete symmetries can be developed.
@@ VI Acknowledgements
We thank Sergey Levine for pointing to relevant references and for discussions on preliminary results, and Hussein Sibai for discussions on symmetry reduction prior to this project.